A body of mass (m) falls from rest through a medium which exact a frictional drag force (bv) proportional to a applied force (f).1. Find it's velocity in terms of time t 2. Find it's terminal velocity
1
Expert's answer
2019-09-09T11:20:04-0400
"ma=m\frac{dv}{dt}=mg-bv+f"

"\frac{dv}{dt}=g-\frac{b}{m}v+\frac{f}{m}"

"bdt=\frac{dv}{\frac{mg}{b}+\frac{f}{b}-v}"

"v(t)=\frac{mg+f}{b}+c\exp{(-bt)}"

"v(0)=0=\frac{mg+f}{b}+c"

So,


"v(t)=\frac{mg+f}{b}(1-\exp{(-bt)})"

Terminal velocity:


"v(\infty)=\frac{mg+f}{b}"